# DiLA: Enhancing LLM Tool Learning with Differential Logic Layer

### v2

@@ 5 Experiments
5.1 Setup
Tasks: We conduct experiments on two fundamental reasoning tasks: the Boolean Satisfiability problem and the Graph Coloring problem. For Boolean Satisfiability problems, we utilize open-source benchmark instances  [^2] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html with 20 to 250 variables, focusing on finding variable assignments that satisfy all constraints. Notably, we do not consider the unsatisfiable cases as finding unsatisfied certificates is a distinct problem. For Graph Coloring problems, we randomly generated 100 3-coloring instances with vertices counts ranging from 10 to 200, following the approach in  Stechly et al. (2024) (see Appendix C for details), aiming to color graph vertices so that no two adjacent vertices share the same color. Furthermore, we also incorporate several complex SAT cases in SAT Competition 2023333https://satcompetition.github.io/2023/ and open-source graph coloring problems with large number of edges [^4] https://www.cs.ubc.ca/~hoos/SATLIB/Benchmarks/SAT/DIMACS/GCP/descr.html to illustrate the robustness of our proposed DiLA. These instances originate from practical industrial problems and frequently pose challenges for modern SAT solvers. Notably, current solvers often enter into an infinite loop when handling these test cases due to their complex structure, requiring a prohibitively long time to solve.
Baselines: We conducted a comparative analysis between DiLA and three baselines, including the prompt-based methods with GPT-4 Achiam et al. (2023) and Llama-3 (Llama-3-70B-Instruct) Meta (2024), and the solver-augmented method, SATLM Ye X (2023). In particular, SATLM Ye X (2023) employs an LLM to parse problem specifications and offloads the logical reasoning task to the Z3 solver De Moura and Bjørner (2008), which serves as the state-of-the-art logic reasoning methods.
Setup: We implement a prototype of our proposed DiLA using Pytorch Paszke et al. (2019), leveraging GPT-4 as the backbone LLM model. Notably, the logic layer within DiLA has no training parameters and can adapt to various problem types expressed in Boolean variables. We employ the Adam optimizer Kingma and Ba (2014) with a learning rate of $2\times 10^{-1}$ in DiLA, facilitating the effective updating of selected variables. Furthermore, we use a temperature of 0.75 for Llama-3 and a higher temperature of 0.9 for GPT-4, consistent with the SATLM approach. We set a time limit of 10,000 seconds for both the solvers and the logic layer.
5.2 Main Results
Table 1: Accuracy and runtime (s) of COT, including GPT-4 and Llama-3, SATLM, and DiLA on simple Boolean Satisfiability and 3-Coloring datasets.
Problem Number of Variables GPT-4 Llama-3 SATLM DiLA
@@ 5.2 Main Results
We report the accuracy of DiLA and the three baselines in Table 1. Accuracy is evaluated based on whether the LLM can output a correct answer that satisfies all constraints. We evaluate LLMs over 100 instances in each domain, with 20 test cases for each scenario. In addition to accuracy, we also report the solving runtime for SATLM and DiLA, both of which leverage GPT-4 as a backbone LLM to perform language understanding. Analysis of Table 1 reveals that both solver-augmented SATLM and logic layer-aided DiLA achieve 100% accuracy when handling SAT problems with up to 250 variables and 3-coloring problems with up to 200 nodes. In contrast, standalone LLMs, including GPT-4 and Llama-3, cannot handle reasoning problems with over 100 variables with COT prompting. More importantly, for all test instances, DiLA exhibits faster performance than SATLM, especially for relatively large cases with over 200 variables. The runtime speedup can be up to 65.5$\times$ when dealing with SAT problems with 250 variables and 23.8$\times$ when tackling 3-coloring problems with 200 nodes. Overall, both the solver-augmented LLM and our proposed DiLA can successfully address these simple artificial test cases, with the runtime speedup highlighting the efficiency of DiLA in logic reasoning.
5.3 Robustness of Reasoning
Compared with SATLM Ye X (2023) and LOGIC-LM Pan et al. (2023), which rely on an off-the-shelf solver for reasoning, DiLA sidesteps the limitations of the solvers themselves, such as high computational costs when dealing with complex formulas. In practice, there are intricate cases where even state-of-the-art SAT solvers struggle, sometimes taking weeks to solve. These challenging instances often involve Boolean satisfiability formulas with complex structures and graph coloring problems with considerable edge connections. For our evaluation, we leverage the state-of-the-art SMT solver, Z3 De Moura and Bjørner (2008), and the widely-used SAT solver, Kissat Biere and Fleury (2022), as baselines, and test on some hard test cases derived from real-world problems.
@@ 6 Related Works
Logic Reasoning in LLMs. Prior approaches to NL-based reasoning with LLMs can be broadly categorized into two groups. One is in-context learning approaches that design special prompts to elicit LLMs’ step-by-step reasoning capabilities. Typical methods include chain-of-thought prompting Wang et al. (2023a) that generates a sequence of reasoning steps before the final answer and the least-to-most prompting Zhou et al. (2022) that breaks the problem down into simpler components that can be solved sequentially. Both the above approaches perform reasoning directly over natural language (NL), providing greater flexibility than symbolic-based reasoning. However, the intrinsic complexity and ambiguity of NL also bring undesired issues such as unfaithful reasoning and hallucinations. The other is tool-augmented approaches that only require LLMs to parse the problem specification out of the language description accurately and leverage off-the-shelf automated tools to derive the final answer. Recent works include SATLM Ye X (2023) and LOGIC-LM Pan et al. (2023), both of which describe the problem with symbolic language using LLMs and then offload the reasoning to external symbolic solvers. The tool-augmented approaches guarantee the correctness of the answer with respect to the parsed specification and avoid planning errors in the solving process. However, the performance of such tool-augmented methods highly relies on external tools and can not deal with many real problems due to the deficiency of backbone solvers.
Neural Symbolic Reasoning. Deep neural networks have made remarkable progress in various domains, but their ability to handle logical reasoning tasks remains uncertain. One potential approach is to encode logical constraints that are essential for certain inference tasks symbolically, making them available to the model either during training or inference. SATNet Wang et al. (2019), for instance, introduces a differentiable MaxSAT solver layer into its network architecture to learn the logical structure of problems through data-driven learning. Their approximation is based on a coordinate descent approach for solving the semidefinite program (SDP) relaxation of the MaxSAT problem. SATNet does not assume that the logical structure of the problem is given, and instead attempts to learn it. In the pursuit of neuro-symbolic learning, SMTLayer Wang et al. (2023b) incorporates a satisfiability modulo theories (SMT) solver into a DNN layer, thereby embedding problem-specific theories into DNN architectures through data-driven training. Compared to SATNet, SMTLayer demonstrates superior accuracy in logical reasoning tasks and a reduced need for training data. However, SMTLayer still requires a certain amount of training data and relies on off-the-shelf SMT solvers for reasoning, making it unsuitable for industrial problem-solving. Similar to the aforementioned neuro-symbolic approach, DiLA also aims to encode logical constraints in the network layer, making them available to the model during forward and backward passes.
@@ 7 Conclusion
In this work, we introduce a pioneering method named differential logic layer-aided language modeling (DiLA). Starting with an NL reasoning problem, DiLA first uses an LLM to cast it into a SAT problem and generate a possible solution based on its language understanding, and then progressively refine this solution within a differential logic layer. In this way, we harness the potential of the language-understanding ability of LLMs and sidestep the limitations of off-the-shelf solvers. Extensive experiments on two reasoning tasks demonstrate the superior efficiency of our approach over state-of-the-art solver-augmented LLMs. Notably, DiLA demonstrates its robustness in dealing with challenging problem instances that SOTA solvers struggle with. The results suggest that DiLA achieves new state-of-the-art performance in symbolic logical reasoning tasks, paving the way for more applications of LLMs in practical reasoning settings.
@@ Appendix A Details of the SAT Specification
To better leverage the parametric knowledge that LLMs have acquired from pretraining on vast amounts of language data, our approach uses a SAT specification as a bridge to encode a range of semantic constraints. Specifically, it translates general reasoning problems into formal constraints with Boolean variables. Below, we provide a prompt example to convert a classical graph 3-coloring problem into its SAT specification.
@@ Appendix B Details of the Initialization
After parsing logic formulas from the problem description, we can directly ask LLM to generate a possible solution according to its language understanding. The possible question prompt can be “What is the logical solution based on the premises?” or “Can you provide a valid answer from these given constraints?”. Here we show a prompt example for a simple boolean satisfiability problem as follows.
@@ Appendix C Dataset Generation
Graph coloring benchmarks belong to the class of NP-complete problems with which exact heuristic-based solvers struggle. To test the effectiveness of our proposed DiLA, we built our dataset using GrinPy [^5] https://pypi.org/project/grinpy/ for common graph operations. Once a successful candidate is found, it is compiled into the standard DIMACS format and appended with a comment containing its precalculated chromatic number. Specifically, before inputting to LLMs, we transform the graph connections in DIMACS format (e.g., 0 3) into natural language statements (e.g., Vertex 0 is connected to vertex 3.). For the experiments in Section 5.2, we generated 100 graph coloring instances with node counts ranging from 10 to 200 and edge counts from 30 to 480. Notably, all these artificial graphs can be successfully colored with 3 colors, providing an ideal and easily solvable setting. An example of one of the graphs with 10 nodes and 30 edges as well as its 3-coloring solution is shown in the following graph.
@@
We also list the semantic edge connections for this 3-coloring problem as follows. After explaining the edge connections given in DIMACS format, LLMs can directly interpret the pure numbers, such as “0 3”, as edge connections, eliminating the need to generate long language sequences for large-scale problems.

### v3

@@ 4 Experiments
Table 1: Accuracy and runtime (s) of CoT, including GPT-4 and Llama-3, LOGIC-LM, and DiLA on LogicalDeduction, simple Boolean Satisfiability and 3-Coloring datasets.
Problem Number of Variables GPT-4 Llama-3 LOGIC-LM DiLA
@@
4.1 Setup
Tasks: We conduct experiments on three fundamental reasoning tasks: the logical deduction problem, Boolean Satisfiability, and Graph Coloring problems. The logic deduction problems are mostly about deducing the order of a sequence of objects from a minimal set of conditions. Here we utilize LogicalDeduction dataset from the BigBench (Srivastava et al., 2022) collaborative benchmark. For Boolean Satisfiability problems, we utilize open-source benchmark instances  [^2] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html with 20 to 250 variables, focusing on finding variable assignments that satisfy all constraints. Notably, we do not consider the unsatisfiable cases as finding unsatisfied certificates is a distinct problem. For Graph Coloring problems, we randomly generated 100 3-coloring instances with vertices counts ranging from 10 to 200, following the approach in  (Stechly et al., 2024) (see Appendix C for details), aiming to color graph vertices so that no two adjacent vertices share the same color. Furthermore, we also incorporate several complex SAT cases in SAT Competition 2023333https://satcompetition.github.io/2023/ and open-source graph coloring problems with large number of edges [^4] https://www.cs.ubc.ca/~hoos/SATLIB/Benchmarks/SAT/ to illustrate the robustness of our proposed DiLA. These instances originate from practical industrial problems and frequently pose challenges for modern SAT solvers. Notably, current solvers often enter into an infinite loop when handling these test cases due to their complex structure, requiring a prohibitively long time to solve.
Baselines: We conducted a comparative analysis between DiLA and three baselines, including the prompt-based methods with GPT-4 (Achiam et al., 2023) and Llama-3 (Llama-3-70B-Instruct) (Meta, 2024), and the solver-augmented method, LOGIC-LM (Pan et al., 2023). In particular, LOGIC-LM (Pan et al., 2023) employs GPT-4 to parse problem specifications and offloads the logical reasoning task to the symbolic solvers, including CSP Solver [^5] https://github.com/python-constraint/ python-constraint and Z3 solver (De Moura and Bjørner, 2008), which serves as the state-of-the-art logic reasoning methods.
Setup: We implement a prototype of our proposed DiLA using Pytorch (Paszke et al., 2019), leveraging GPT-4 as the backbone LLM model. Notably, the logic layer within DiLA has no training parameters and can adapt to various problem types expressed in Boolean variables. We employ the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $2\times 10^{-1}$ in DiLA, facilitating the effective updating of selected variables. Furthermore, we use a temperature of 0 for LLMs, consistent with the LOGIC-LM approach. We set a time limit of 10,000 seconds for both the solvers and the logic layer.
4.2 Main Results
@@
We report the accuracy of DiLA and the three baselines in Table 1. Accuracy is evaluated based on whether the LLM can output a correct answer that satisfies all constraints. We evaluate LLMs over 100 instances in each domain. In addition to accuracy, we also report the solving runtime for LOGIC-LM and DiLA, both of which leverage GPT-4 as a backbone LLM to perform language understanding. Analysis of Table 1 reveals that both solver-augmented LOGIC-LM and logic layer-aided DiLA achieve 100% accuracy on these simple benchmarks, owing to the precise logic parsing and accurate solving. In contrast, standalone LLMs, including GPT-4 and Llama-3, cannot handle reasoning problems with over 100 variables with CoT prompting. More importantly, for all test instances, DiLA exhibits faster performance than LOGIC-LM in solving process, especially for relatively large cases with over 200 variables. The runtime speedup can be up to 65.5$\times$ when dealing with SAT problems with 250 variables and 23.8$\times$ when tackling 3-coloring problems with 200 nodes. Overall, both the solver-augmented LLM and our proposed DiLA can successfully address these simple artificial test cases, with the runtime speedup highlighting the efficiency of DiLA in logic reasoning.
@@
4.3 Robustness of Reasoning
Compared with SATLM (Ye X, 2023) and LOGIC-LM (Pan et al., 2023), which rely on an off-the-shelf solver for reasoning, DiLA sidesteps the limitations of the solvers themselves, such as high computational costs when dealing with complex formulas. In practice, there are intricate cases where even state-of-the-art SAT solvers struggle, sometimes taking weeks to solve. These challenging instances often involve Boolean satisfiability formulas with complex structures and graph coloring problems with considerable edge connections. For our evaluation, we leverage the state-of-the-art SMT solver, Z3 (De Moura and Bjørner, 2008), and the widely-used SAT solver, Kissat (Biere and Fleury, 2022), as baselines, and test on some hard test cases derived from real-world problems.
@@ 5 Related Works
Prior approaches to NL-based reasoning with LLMs can be broadly categorized into two groups. One is in-context learning approaches that design special prompts to elicit LLMs’ step-by-step reasoning capabilities. Typical methods include chain-of-thought prompting (Wang et al., 2023) that generates a sequence of reasoning steps before the final answer and the least-to-most prompting (Zhou et al., 2022) that breaks the problem down into simpler components that can be solved sequentially. Both the above approaches perform reasoning directly over natural language (NL), providing greater flexibility than symbolic-based reasoning. However, the intrinsic complexity and ambiguity of NL also bring undesired issues such as unfaithful reasoning and hallucinations. The other is tool-augmented approaches that only require LLMs to parse the problem specification out of the language description accurately and leverage off-the-shelf automated tools to derive the final answer, as exemplified by SATLM (Ye X, 2023) and LOGIC-LM (Pan et al., 2023). The tool-augmented approaches guarantee the correctness of the answer with respect to the parsed specification and avoid planning errors in the solving process. However, the performance of such tool-augmented methods highly relies on external tools and can not deal with many real problems due to the deficiency of backbone solvers.
@@ 6 Conclusion
In this work, we introduce a pioneering method named differential logic layer-aided language modeling (DiLA). Starting with an NL reasoning problem, DiLA first uses an LLM to cast it into a SAT problem and generate a possible solution based on its language understanding, and then progressively refines this solution within a logic layer. In this way, we harness the potential of the language-understanding ability of LLMs and sidestep the limitations of off-the-shelf solvers. Extensive experiments on two reasoning tasks demonstrate the superior efficiency of our approach over state-of-the-art solver-augmented LLMs. Notably, DiLA demonstrates its robustness in dealing with challenging problem instances that SOTA solvers struggle with. The results suggest that DiLA achieves new state-of-the-art performance in symbolic logical reasoning tasks, paving the way for more applications of LLMs in practical reasoning settings.
7 Limitation
We identify two main limitations of DiLA. First, DiLA relies on translating reasoning problems into logical formats that can be encoded by boolean variables and first-order logic. As a consequence, the model’s applicability is inherently bounded by the expressiveness of the underlying theory, as not all problems can be easily encoded in first-order formulas with boolean vectors. Nevertheless, the wide range of tasks that we can instantiate our DiLA framework on shows its general applicability.
Second, DiLA relies on backbone LLMs to translate natural language problems into symbolic representations. Although this approach has shown promise, it may struggle with logical representations featuring complex grammar structures, such as probabilistic circuit logic. This limitation stems from the challenge of communicating intricate grammatical rules to the language model through a limited number of demonstrations within a constrained context size. To address this, future research could investigate the development of specialized modules to improve the mapping between natural language and symbolic language, for instance, by fine-tuning LLMs with symbolic formulas for specific reasoning tasks.
@@ Appendix A Details of the SAT Specification
To better leverage the parametric knowledge that LLMs have acquired from pretraining on vast amounts of language data, our approach uses a SAT specification as a bridge to encode a range of semantic constraints. Specifically, it translates general reasoning problems into formal constraints with Boolean variables. Below, we provide a prompt example of how to convert a classical graph 3-coloring problem into its SAT specification in Table 3.
@@
Table 3: Prompt example of converting a classical graph 3-coloring problem into its SAT specification.
@@ Appendix B Details of the Initialization
After parsing logic formulas from the problem description, we can directly ask LLM to generate a possible solution according to its language understanding. The possible question prompt can be “What is the logical solution based on the premises?” or “Can you provide a valid answer from these given constraints?”. Here we show a prompt example for a simple boolean satisfiability problem in Table 4.
@@ 2. A = false, B = true, C = true
Table 4: Example of generating an initial solution for a boolean satisfiability problem.
@@ Appendix C Dataset Generation
Graph coloring benchmarks belong to the class of NP-complete problems with which exact heuristic-based solvers struggle. To test the effectiveness of our proposed DiLA, we built our dataset using GrinPy [^6] https://pypi.org/project/grinpy/ for common graph operations. Once a successful candidate is found, it is compiled into the standard DIMACS format and appended with a comment containing its precalculated chromatic number. Specifically, before inputting to LLMs, we transform the graph connections in DIMACS format (e.g., 0 3) into natural language statements (e.g., Vertex 0 is connected to vertex 3.). For the experiments in Section 4.2, we generated 100 graph coloring instances with node counts ranging from 10 to 200 and edge counts from 30 to 480. Notably, all these artificial graphs can be successfully colored with 3 colors, providing an ideal and easily solvable setting. An example of one of the graphs with 10 nodes and 30 edges as well as its 3-coloring solution is shown in Figure 5.
@@
Figure 5: 3-coloring graph example.
We also list the semantic edge connections for this 3-coloring problem in Table 5. After explaining the edge connections given in DIMACS format, LLMs can directly interpret the pure numbers, such as “0 3”, as edge connections, eliminating the need to generate long language sequences for large-scale problems.
@@
Table 5: Edge connections in language for a 3-coloring problem.